OpenStudy (jjwalton04):

Simplify square root of 3 multiplied by the fifth root of 3. please EXPLAIN

OpenStudy (jjwalton04):

\[Simplfy \left( \sqrt{3} \right)\left( ^5\sqrt{3} \right)\]

OpenStudy (triciaal):